# Fiedler-Based Characterization and Identification of Leaders in Semi-Autonomous Networks

### v1

@@ Abstract
This paper addresses the problem of identifying leader nodes in semi-autonomous consensus networks from observed agent dynamics. Using the grounded Laplacian formulation, we derive spectral conditions that ensure the components of the Fiedler vector associated with leader and follower nodes are distinct. Building on this foundation, we employ the notion of relative tempo from prior work as an observable quantity that relates agents’ steady-state velocities to the Fiedler vector. This relationship enables the development of a data-driven algorithm that reconstructs the Fiedler vector—and consequently identifies the leader set—using only steady-state velocity measurements, without requiring knowledge of the network topology. The proposed approach is validated through numerical examples, demonstrating how spectral properties and relative tempo measurements can be combined to reveal hidden leadership structures in consensus networks.
@@ I INTRODUCTION
The ability to infer or influence the structure of a networked dynamical system is central to understanding and controlling collective behaviors in multi-agent systems. In many practical scenarios, agents interact through local rules without global supervision, and only partial measurements of their states are available. Identifying how such interactions shape the network dynamics has therefore become an important research direction in network systems theory [1, 2, 3, 4, 5].
The network identification problem concerns the task of reconstructing or inferring the underlying interaction topology of a multi-agent system from observed data. In other words, given measurements of the agents’ states or outputs, one seeks to determine which agents exchange information and with what strength. This problem is fundamental in areas such as systems biology [6], neuroscience [7], social networks [8], and engineering systems [9]. A central challenge is that the observed collective behavior does not uniquely determine the network structure without additional assumptions or prior knowledge. Accordingly, approaches to network identification are typically organized along a spectrum: at one end, methods assume complete dynamical models and seek to refine unknown network parameters, while at the other end, data-driven methods attempt to infer both the dynamics and the topology simultaneously [10, 11]. In the context of multi-agent coordination, identifying the interaction graph is crucial for understanding how local rules give rise to global behaviors, as well as for designing control strategies that are robust to uncertainties or malicious attacks in the communication topology. An illustration of network topology identification is presented in Figure 1.
@@
Early work focused on offline reconstruction of the interaction topology or coupling weights from state or input–output data, often assuming access to global trajectories [9]. More recent efforts have sought weaker conditions for identifiability, studying local structural properties, spectral methods, or algebraic constraints that ensure unique reconstruction up to graph symmetries [12, 13]. Other works approach the problem by probing the network with specially designed inputs, [14, 15], or by node knock-out [16].
In many real-world networks, not all agents behave identically—some may possess additional autonomy or receive external commands. These semi-autonomous consensus networks consist of a subset of agents, often termed leaders or externally driven nodes, that are subject to independent control inputs or reference signals, while the remaining followers update their states through local interactions within the network [17, 18, 19]. Such leader–follower architectures appear in diverse applications including vehicle platooning, robotic swarms, and opinion dynamics, where overall coordination must be preserved despite partial autonomy.
From a systems-theoretic viewpoint, these networks introduce an additional layer of structure atop the standard consensus model: the interconnection topology not only determines how information flows, but also which nodes actively influence collective behavior. Consequently, an important question is how to identify the leaders—that is, determine which agents receive external inputs or exert disproportionate influence on the group—using only observed data. This leader identification problem can be regarded as a focused instance of the broader network identification problem, where the goal is to infer the underlying interaction structure or influence pattern from measurements of the agents’ trajectories [20]. To the best of our knowledge, this specific problem has received little direct attention in the literature.
The main contribution of this paper is the development of a data-driven algorithm for identifying leader nodes in semi-autonomous consensus networks. Building on a new spectral characterization of leader–follower graphs, we first establish analytical conditions under which the components of the Fiedler vector corresponding to leaders and followers are separable. We then leverage the concept of relative tempo from [21] as a measurable proxy for the Fiedler vector, enabling estimation of the network’s influence structure directly from steady-state velocity data. This connection leads to a simple yet effective leader identification procedure that does not require knowledge of the underlying topology. The proposed framework therefore unifies spectral graph analysis with data-driven inference and provides new insights into how leader influence manifests in the dynamical response of the network.
@@
Let $\mathbb{R}$ and $\mathbb{N}$ denote the set of real numbers and natural numbers, respectively, and $\mathbb{R}^{m\times n}$ the space of real matrices of size $m\times n$. The $n$-dimensional vector of all ones is denoted $\mathbbm{1}_{n}$, and $I_{n}$ is the identity matrix. A graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W})$ is defined by a set of vertices $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, and weights $\mathcal{W}:\mathcal{E}\to\mathbb{R}$. For directed edges, $e_{ij}=(v_{i},v_{j})\in\mathcal{E}$ denotes a directed edge from $v_{i}$ to $v_{j}$, while for undirected edges $e_{ij}=\{v_{i},v_{j}\}$. The notation $i\sim j$ indicates that an edge exists between the nodes $i$ and $j$, while $\mathcal{N}(i)=\{j\in\mathcal{V}\,|\,i\sim j\}$ is the neighbor set of node $i$. The degree of node $i$, $\deg(i)$, is the cardinality of its neighbor set.
@@ II PROBLEM STATEMENT
We now introduce a basic assumption about the number of leaders and followers. It states that there must be at least one follower node, and no leader node should be connected to all the other nodes.
@@
The follower set is non-empty ($\mathcal{V}_{f}\neq\emptyset$). Furthermore, $\deg(i)<|\mathcal{V}|-1$ for all $v_{i}\in\mathcal{V}_{\ell}$.
Each agent implements the celebrated consensus protocol [22], modified such that the leader agents receive the additional input $u_{i}^{ex}$. That is, each agent has the closed-loop dynamics
@@
We now observe that the matrix $\bar{L}_{11}$ in (3) is the grounded Laplacian of the undirected version of $\mathcal{\bar{G}}$ [23], where the external control nodes (the $y_{i}$’s) are removed. The Fiedler eigenvalue, $\lambda_{F}$ corresponds to the smallest eigenvalue of $\bar{L}_{11}$; its corresponding eivenvector $v_{F}$ can be chosen to have all positive entries. It was shown in [23] that under Assumption 1 that $0<\lambda_{F}<1$. For the remainder of this work, we refer to the Fiedler eigenvector of $\mathcal{G}$ to mean the Fiedler vector of the grounded Laplacian $\bar{L}_{11}$.
@@ III Leader-Identifiable Graphs
It is well-known in the literature that the Fiedler eigenvector can be used to identify clusters in a graph [24, 25]. In this section we derive conditions that ensure the elements of the Fiedler vector can be separated into two groups, one corresponding to the follower nodes and one to the leader nodes.
We now explore algebraic and combinatorial connections between the Fiedler vector associated with $\bar{L}_{11}$, and the structure of the graph $\mathcal{G}$. To begin, we first observe that the entries of the Fiedler vector can be related to walks on an associated normalized graphs. In this direction, we define the semi-normalized adjacency matrix of $\mathcal{G}$ as
@@
Because $\hat{A}$ is a nonnegative matrix whose associated graph is strongly connected, the Perron–Frobenius theorem [26] implies that its spectral radius is real, positive, and simple, and that the corresponding eigenvector can be chosen strictly positive. Since $v_{F}$ has positive entries, the pair $(\lambda,v_{F})=(1,v_{F})$ is the unique Perron–Frobenius eigenpair of $\hat{A}(\mathcal{G})$. ∎
@@
Furthermore, let $\underline{\deg}_{f}^{\sigma(i)}=\min_{j\in\mathcal{V}_{f}^{\sigma(i)}}\deg_{f}^{\sigma(i)}(j)$ denote the minimum follower degree. The sequence $\{\mathcal{G}^{\sigma(i)}\}$ satisfies the following properties:
@@ IV A Leader Identification Algorithm
We next review the notion of relative tempo introduced in [21], which extends the earlier concept of the degree of relative influence proposed in [27]. These metrics quantify how the evolution rate of one agent (or subgroup) compares to another within a consensus-type network. Formally, the relative tempo between two agents is defined as
@@
For large times $t>T$, the agents’ dynamics are dominated by the slowest decaying mode of the network Laplacian. Projecting the system onto its eigenbasis and neglecting the zero and fast-decaying modes yields [27]
@@
We are now prepared to propose an algorithm to identify the leaders in a semi-autonomous consensus network with graphs satisfying the properties of Definition 1 and Theorem 1.

### v2

@@ Abstract
This paper addresses the problem of identifying leader nodes in semi-autonomous consensus networks from observed agent dynamics. Using the grounded Laplacian formulation, we derive graph-theoretic conditions that ensure the components of the Fiedler vector associated with leader and follower nodes are distinct. Building on this foundation, we employ the notion of relative tempo from prior work as an observable quantity that relates agents’ steady-state velocities to the Fiedler vector. This relationship enables the development of a data-driven algorithm that reconstructs the Fiedler vector—and consequently identifies the leader set—using only steady-state velocity measurements, without requiring knowledge of the network topology. The proposed approach is validated through numerical examples, demonstrating how graph-theoretic properties and relative tempo measurements can be combined to reveal hidden leadership structures in consensus networks.
@@ I INTRODUCTION
The ability to infer or influence the structure of a networked dynamical system is central to understanding and controlling collective behaviors in multi-agent systems. In many practical scenarios, agents interact through local rules without global supervision, and only partial measurements of their states are available. Identifying how such interactions shape the network dynamics has therefore become an important research direction in network systems theory [Goncalves2008, Shahrampour2013, Hendrickx2018, GEVERS201710580, WEERTS2018247].
The network identification problem concerns the task of reconstructing or inferring the underlying interaction topology of a multi-agent system from observed data. In other words, given measurements of the agents’ states or outputs, one seeks to determine which agents exchange information and with what strength. This problem is fundamental in areas such as systems biology [Tegner2003], neuroscience [Friston2011], social networks [Easley2010], and engineering systems [Materassi2012]. A central challenge is that the observed collective behavior does not uniquely determine the network structure without additional assumptions or prior knowledge. Accordingly, approaches to network identification are typically organized along a spectrum: at one end, methods assume complete dynamical models and seek to refine unknown network parameters, while at the other end, data-driven methods attempt to infer both the dynamics and the topology simultaneously [Timme2014, Ren2007]. In the context of multi-agent coordination, identifying the interaction graph is crucial for understanding how local rules give rise to global behaviors, as well as for designing control strategies that are robust to uncertainties or malicious attacks in the communication topology. An illustration of network topology identification is presented in Figure 1.
@@
Early work focused on offline reconstruction of the interaction topology or coupling weights from state or input–output data, often assuming access to global trajectories [Materassi2012]. More recent efforts have sought weaker conditions for identifiability, studying local structural properties, graph-theoretic methods, or algebraic constraints that ensure unique reconstruction up to graph symmetries [VANWAARDE2018319, 8382192]. Other works approach the problem by probing the network with specially designed inputs, [Timme2007, NetworkIdentificationDiffusivelyCoupledNetworksMinimalTimeComplexity], or by node knock-out [Nabi2010].
In many real-world networks, not all agents behave identically—some may possess additional autonomy or receive external commands. These semi-autonomous consensus networks consist of a subset of agents, often termed leaders or externally driven nodes, that are subject to independent control inputs or reference signals, while the remaining followers update their states through local interactions within the network [NI2010209, REN2007474, Li01012019]. Such leader–follower architectures appear in diverse applications including vehicle platooning, robotic swarms, and opinion dynamics, where overall coordination must be preserved despite partial autonomy.
From a systems-theoretic viewpoint, these networks introduce an additional layer of structure atop the standard consensus model: the interconnection topology not only determines how information flows, but also which nodes actively influence collective behavior. Consequently, an important question is how to identify the leaders—that is, determine which agents receive external inputs or exert disproportionate influence on the group—using only observed data. This leader identification problem can be regarded as a focused instance of the broader network identification problem, where the goal is to infer the underlying interaction structure or influence pattern from measurements of the agents’ trajectories [Dai2019]. To the best of our knowledge, this specific problem has received little direct attention in the literature.
The main contribution of this paper is the development of a data-driven algorithm for identifying leader nodes in semi-autonomous consensus networks. Building on a new spectral characterization of leader–follower graphs, we first establish analytical conditions under which the components of the Fiedler vector corresponding to leaders and followers are separable. We then leverage the concept of relative tempo from [Shao2019relative] as a measurable proxy for the Fiedler vector, enabling estimation of the network’s influence structure directly from steady-state velocity data. This connection leads to a simple yet effective leader identification procedure that does not require knowledge of the underlying topology. The proposed framework therefore unifies graph analysis with data-driven inference and provides new insights into how leader influence manifests in the dynamical response of the network.
@@
Let $\mathbb{R}$ and $\mathbb{N}$ denote the set of real numbers and natural numbers, respectively, and $\mathbb{R}^{m\times n}$ the space of real matrices of size $m\times n$. The $n$-dimensional vector of all ones is denoted $\mathds{1}_{n}$, and $I_{n}$ is the identity matrix. A graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W})$ is defined by a set of vertices $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, and weights $\mathcal{W}:\mathcal{E}\to\mathbb{R}$. For directed edges, $e_{ij}=(v_{i},v_{j})\in\mathcal{E}$ denotes a directed edge from $v_{i}$ to $v_{j}$, while for undirected edges $e_{ij}=\{v_{i},v_{j}\}$. The notation $i\sim j$ indicates that an edge exists between the nodes $i$ and $j$, while $\mathcal{N}(i)=\{j\in\mathcal{V}\,|\,i\sim j\}$ is the neighbor set of node $i$. The degree of node $i$, $\deg(i)$, is the cardinality of its neighbor set.
@@ II PROBLEM STATEMENT
We now introduce a basic assumption about the number of leaders and followers. It states that there must be at least one follower node.
@@
The follower set is non-empty ($\mathcal{V}_{f}\neq\emptyset$).
Each agent implements the celebrated consensus protocol [Mesbahi2010], modified such that the leader agents receive the additional input $u_{i}^{ex}$. That is, each agent has the closed-loop dynamics
@@
We now observe that the matrix $\bar{L}_{11}$ in (3) is the grounded Laplacian of the undirected version of $\mathcal{\bar{G}}$ [OnTheSmallestEigenvectorGroundedLaplac], where the external control nodes (the $y_{i}$’s) are removed. The Fiedler eigenvalue, $\lambda_{F}$ corresponds to the smallest eigenvalue of $\bar{L}_{11}$; its corresponding eivenvector $v_{F}$ can be chosen to have all positive entries. It was shown in [OnTheSmallestEigenvectorGroundedLaplac] that under Assumption 1 that $0<\lambda_{F}<1$. For the remainder of this work, we refer to the Fiedler eigenvector of $\mathcal{G}$ to mean the Fiedler vector of the grounded Laplacian $\bar{L}_{11}$.
@@ III Leader-Identifiable Graphs
It is well-known in the literature that the Fiedler eigenvector can be used to identify clusters in a graph [Ng_NIPS21, Luxburg2007]. In this section we derive conditions that ensure the elements of the Fiedler vector can be separated into two groups, one corresponding to the follower nodes and one to the leader nodes.
We now explore algebraic and combinatorial connections between the Fiedler vector associated with $\bar{L}_{11}$, and the structure of the graph $\mathcal{G}$. In this direction, we define the semi-normalized adjacency matrix of $\mathcal{G}$ as
@@
Because $\hat{A}$ is a nonnegative matrix whose associated graph is strongly connected, the Perron–Frobenius theorem [horn13] implies that its spectral radius is real, positive, and simple, and that the corresponding eigenvector can be chosen strictly positive. Since $v_{F}$ has positive entries, the pair $(\lambda,v_{F})=(1,v_{F})$ is the unique Perron–Frobenius eigenpair of $\hat{A}(\mathcal{G})$. ∎
@@
Furthermore, let $\underline{\deg}_{f}^{\sigma(i)}=\min_{j\in\mathcal{V}_{f}^{\sigma(i)}}\deg_{f}^{\sigma(i)}(j)$ denote the minimum follower-follower degree The sequence $\{\mathcal{G}^{\sigma(i)}\}$ satisfies the following properties:
@@ IV A Leader Identification Algorithm
We next review the notion of relative tempo introduced in [Shao2019relative], which extends the earlier concept of the degree of relative influence proposed in [Mesbahi2014]. These metrics quantify how the evolution rate of one agent (or subgroup) compares to another within a consensus-type network. Formally, the relative tempo between two agents is defined as
@@
For large times $t>T$, the agents’ dynamics are dominated by the slowest decaying mode of the network Laplacian. Projecting the system onto its eigenbasis and neglecting the zero and fast-decaying modes yields [Mesbahi2014]
@@
We now propose an algorithm to identify leaders in semi-autonomous consensus networks whose graphs satisfy Definition 1 and Theorem 1.